OpenStudy (anonymous):

help me please so badly need help on this √9:(x-7)*√9:((x-7))^(7)

OpenStudy (anonymous):

i will help if i can read it

OpenStudy (anonymous):

i think 9/ (x-7)^8

OpenStudy (anonymous):

\[\sqrt{9}(x-7)\sqrt{9}(x-7)^7=9(x-7)^8\]\]

OpenStudy (anonymous):

suzi you must have read what i did. wasn't sure

OpenStudy (anonymous):

i thought the question is √(9 / (x-7)) * √(9 / ((x-7))^(7)) the answer would be 9 / (x-7)^8 :D